A 200-W infrared laser emits photons with a wavelength of 8.60 10-7 m, and a 200-W ultraviolet light emits photons with a wavelength of 2.64 10-7 m.

(a) Which has greater energy, a single infrared photon or a single ultraviolet photon?

-a single infrared photon


-a single ultraviolet photon
(b) What is the energy of a single infrared photon?

What is the energy of a single ultraviolet photon?

(c) How many infrared photons are emitted per second?

How many ultraviolet photons are emitted per second?
Answer #1

Part a & b:

Energy of single photon emitted from infrared laser light,

E_{laser}=\frac{hc}{\lambda }=\frac{6.62\times 10^{-34}\times 3\times 10^{8}}{8.60\times 10^{-7}}=2.3\times 10^{-19}J

Energy of single photon emitted from ultraviolet light,

E_{uv}=\frac{hc}{\lambda }=\frac{6.62\times 10^{-34}\times 3\times 10^{8}}{2.64\times 10^{-7}}=7.52\times 10^{-19}J

So the single ultraviolet photon has greater energy.

Part C:

No . of photons emitted/see from the infrared source,

N_{infrared}=\frac{200}{2.3\times 10^{-19}}=8.69\times 10^{20}

No . of photons emitted/see from the UV source,

N_{uv}=\frac{200}{7.52\times 10^{-19}}=2.65\times 10^{20}
Answer #2

a)

b) E = hc/? = 7.42 * 10^-18 J forUV

   E = 2.3 * 10^-18 for infra